# Learning Noise-Robust Stable Koopman Operator for Control with Physics-Informed Observables

### v1

@@ III-B Control Performance
The trained Koopman model is used as a surrogate model for model predictive control, following the approach in [15]. The prediction horizon, $N_{P}$, is set to 20 steps. Control input constraints are defined as $u_{\text{min}}=-20$ and $u_{\text{max}}=20$, with no constraints on the system states. For the cost function described in (17), $Q_{N_{P}}$ and $Q_{N}$ are set to ones on the first $n$ diagonal elements, with zeros elsewhere. The parameter $R$ is set to 0.
@@ Results
Figure 5 compares the Model Predictive Control (MPC) costs for the CartPole system using three different models under a noise level of 0.100. The models compared are EDMD with RBF basis functions (eDMD-RBF), the Dissipative Koopman model with Polyflow basis functions (Dis-PF), and the Standard Koopman model with Polyflow basis functions (Std-PF). The violin plots illustrate the distribution of MPC costs across various initial states. The results indicate that both the Dissipative Koopman model with Polyflow basis functions (Dis-PF) and the Standard Koopman model with Polyflow basis functions (Std-PF) achieve relatively similar MPC costs. Both models show stable control performance across different initial states, demonstrating robustness in control. In contrast, the EDMD with RBF basis functions (eDMD-RBF) exhibits higher variability in MPC costs and generally incurs higher average MPC costs compared to the Dissipative and Standard Koopman models. This suggests that while the EDMD-RBF model provides some level of stability, it does so at the expense of higher control costs and less consistency. Overall, the comparison reveals that both the Dissipative and Standard Koopman models with Polyflow basis functions offer superior and more stable control performance compared to the EDMD-RBF model under the noise level of 0.1000. This highlights the effectiveness of the Polyflow basis functions in achieving robust control in MPC applications.
@@ IV Iterative Data Augmentation
Figure 6 visualizes the trajectory of the CartPole system across three different scenarios: no data augmentation, 1x data augmentation, and 2x data augmentation. In the first scenario, where no data augmentation was applied, the Koopman model fails to stabilize the CartPole after 200 steps. However, when the failed trajectory, along with other failed trajectories from different initial conditions, is added to the training data, the Koopman model demonstrates significantly improved control outcomes. The system’s performance notably improves with data augmentation, transitioning from oscillating to stable trajectories. This enhancement is also reflected in the reduction of the Model Predictive Control (MPC) cost, with costs decreasing from 135.6 (no augmentation) to 74.7 (1x augmentation) and 67.8 (2x augmentation).
@@
Figure 7 presents violin plots illustrating the distribution of Model Predictive Control (MPC) costs for the CartPole system across two different tasks—seen and unseen—and iterative models under a noise level of 0.1000. The “seen” task refers to scenarios where input bounds are set between $[-20,20]$, while the “unseen” task refers to tighter input bounds of $[-10,10]$. Data augmentation was performed using closed loop trajectories from the seen task. The violin plots display MPC costs for four different initial states and two levels of data augmentation, comparing the costs for both control tasks. Each subplot emphasizes the performance gains achieved through iterative learning. The results indicate that both seen and unseen control tasks benefit significantly from iterative learning, with successive iterations leading to lower and more stable MPC costs. This demonstrates the effectiveness of iterative learning in refining control strategies and enhancing the overall performance of the system.
@@ V Ablation Studies
In this section, we will analyze the individual effect of the following components within our framework on the overall model performance with training data at different noise levels:
@@ V-A Dissipative Koopman Model versus Standard Koopman Model
Figure 9(a) compares the performance of the Dissipative Koopman Model and the Standard Koopman Model across different noise levels for multiple trajectories. The Dissipative Koopman Model consistently shows lower mean errors and mean test errors, underscoring its superior prediction accuracy and robustness. However, the training time for the Dissipative Koopman Model is considerably longer than that of the Standard Koopman Model, due to the increased number of learning parameters and the computationally intensive operations, such as matrix exponentiation and pseudo-inverse computation. Despite its relatively short training time, a significant drawback of the Standard Koopman Model is its sensitivity to the initial conditions of the Koopman matrix. If these initial conditions are not carefully tuned, the model can produce unstable eigenvalues, leading to poor performance. Figure 8 illustrates this issue by showing the discrete-time eigenvalues of the standard and dissipative Koopman operators for the Van Der Pol Oscillator System at a noise level of 0.0599, using a different weight initialization than in previous cases. For fair comparison, both standard and dissipative Koopman models are initialized with stable Koopman matrices. However, after 1,000 training epochs, some of the discrete-time eigenvalues in the Standard Koopman matrix move outside the unit circle, leading to a loss of stability while dissipative Koopman model never loses stability during training.
@@ V-B Learning Rate Scheduler
Figure 9(b) compares the performance of the model using constant learning rate (LR) scheduling and cyclic learning rate scheduling under different noise levels. The cyclic learning rate scheduling generally results in slightly lower mean errors and mean test errors compared to the constant learning rate. The time taken for training with cyclic learning rate scheduling is slightly higher due to the oscillation in learning rates.
@@ V-C Optimizer Switching
Figure 9(c) examines the effect of switching optimizers on the model’s performance. Different combinations of Adam and LBFGS optimizers were used, and their impact on mean errors, mean test errors, and training times was analyzed. Switching optimizers can lead to lower mean errors and test errors in some configurations, suggesting that optimizer switching can enhance training effectiveness. The combination of Adam and LBFGS every 500 epochs generally shows better performance with lower errors but increases the training time significantly. The mean error and test error for configurations where optimizers are switched show improved stability and reduced overfitting.
V-D Progressive roll-out Length Vs. Constant roll-out Length
Figure 9(d) compares the performance of the model using progressive roll-out length scheduling and constant roll-out length scheduling under different noise levels. The progressive roll-out scheduling shows lower mean errors and mean test errors compared to the constant roll-out length, indicating better learning and generalization. The time taken for training with progressive roll-out scheduling is higher due to the incremental increase in roll-out length.
@@
Figure 9: Ablation study on the Van Der Pol Oscillator system. First column: training error. Second column: testing error. Third column: computational wall time.
@@ VI Conclusions
In this work, we proposed a novel data-driven learning framework for the Koopman operator of nonlinear dynamical systems. Unlike existing frameworks that rely on ad-hoc observables or complex neural networks to approximate the Koopman invariant subspace, our approach utilizes observables chosen from known system dynamics with closure through the concept of Polyflow. This physics-informed basis significantly enhances the performance and robustness of the resulting models compared to ad-hoc choices. To ensure noise robustness and long-term stability of the lifted system, we designed a stable parameterization of the Koopman operator and employed a roll-out recurrent loss. Our stabilized parameterization guarantees a stable system, which can be equally accurate or, in some cases, even more accurate than the standard Koopman operator. Additionally, we addressed the curse of dimensionality in sampling from high-dimensional phase space by iteratively augmenting data from closed-loop trajectories. This approach not only improved performance in seen closed-loop scenarios but also enhanced the model’s ability to handle unseen control objectives effectively. We conducted an ablation study to demonstrate the effectiveness of several key components: switching to the LBFGS optimizer for fine-tuning, employing cyclic learning rate scheduling, utilizing progressive roll-out lengths for enhanced training outcomes, and incorporating a stabilized Koopman operator to ensure guaranteed system stability. Our numerical results on the Van der Pol Oscillator and the CartPole system demonstrated the superior performance of the proposed framework in prediction and model predictive control (MPC) of nonlinear systems, compared to other popular DMD-based methods Overall, our framework offers a noise-robust approach for learning stable Koopman operators, providing a solid foundation for future research and applications in nonlinear dynamical system control.

### v2

@@ III-B Control Performance
The trained Koopman model is used as a surrogate model for model predictive control, following the approach in [18]. The prediction horizon, $N_{P}$, is set to 20 steps. Control input constraints are defined as $u_{\text{min}}=-20$ and $u_{\text{max}}=20$, with no constraints on the system states. For the cost function described in eq. 17, $Q_{N_{P}}$ and $Q_{N}$ are set to ones on the first $n$ diagonal elements, with zeros elsewhere. The parameter $R$ is set to 0.
@@ Results
Figure 5 compares the Model Predictive Control (MPC) costs for the CartPole system using three different models under a noise level of 0.100. The models compared are EDMD with RBF basis functions (eDMD-RBF), the Dissipative Koopman model with Polyflow basis functions (Dis-PF), and the Standard Koopman model with Polyflow basis functions (Std-PF). The violin plots illustrate the distribution of MPC costs across various initial states. The results indicate that both the Dissipative Koopman model with Polyflow basis functions (Dis-PF) and the Standard Koopman model with Polyflow basis functions (Std-PF) achieve relatively similar MPC costs. Both models show stable control performance across different initial states, demonstrating robustness in control. In contrast, the EDMD with RBF basis functions (eDMD-RBF) exhibits higher variability in MPC costs and generally incurs higher average MPC costs compared to the Dissipative and Standard Koopman models. This suggests that while the EDMD-RBF model provides some level of stability, it does so at the expense of higher control costs and less consistency. Overall, the comparison reveals that both the Dissipative and Standard Koopman models with Polyflow basis functions offer superior and more stable control performance compared to the EDMD-RBF model under the noise level of 0.1. This highlights the effectiveness of the Polyflow basis functions in achieving robust control in MPC applications.
@@ IV Iterative Data Augmentation
Figure 6 visualizes the trajectory of the CartPole system across three different scenarios: no data augmentation, one-time data augmentation, and twice data augmentation. In the first scenario, where no data augmentation was applied, the Koopman model fails to stabilize the CartPole after 200 steps. However, when the failed trajectory, along with other failed trajectories from different initial conditions, is added to the training data, the Koopman model demonstrates significantly improved control outcomes. The performance notably improves with data augmentation, transitioning from oscillating to stable trajectories. This enhancement is also reflected in the reduction of the Model Predictive Control (MPC) cost, with costs decreasing from 135.6 (no augmentation) to 74.7 (one-time augmentation) and 67.8 (twice augmentation).
@@
Figure 7 presents violin plots illustrating the distribution of Model Predictive Control (MPC) costs for the CartPole system across two different tasks—seen and unseen—and iterative models under a noise level of 0.1. The “seen” task refers to scenarios where input bounds are set between $[-20,20]$, while the “unseen” task refers to tighter input bounds of $[-10,10]$. Data augmentation was performed using closed loop trajectories from the seen task. It is important to note that classical iterative learning [50] cannot easily generalize to a different task. The violin plots display MPC costs for four different initial states and two levels of data augmentation, comparing the costs for both control tasks. Each subplot emphasizes the performance gains achieved through iterative learning. The results indicate that both seen and unseen control tasks benefit significantly from iterative learning, with successive iterations leading to lower and more stable MPC costs. This demonstrates the effectiveness of iterative learning in enhancing the overall performance of the system.
@@ V Ablation Studies
In this section, we will analyze the individual effect of the following components on the overall model performance with training data at different noise levels:
@@ V-A Dissipative Koopman Model versus Standard Koopman Model
Figure 9a compares the performance of the Dissipative Koopman Model and the Standard Koopman Model across different noise levels for multiple trajectories. The Dissipative Koopman Model consistently shows lower mean errors and mean test errors, underscoring its superior prediction accuracy and robustness. However, the training time for the Dissipative Koopman Model is considerably longer than that of the Standard Koopman Model, due to the increased number of learning parameters and the computationally intensive operations, such as matrix exponentiation and pseudo-inverse computation. Despite its relatively short training time, a significant drawback of the Standard Koopman Model is its sensitivity to the initial conditions of the Koopman matrix. If these initial conditions are not carefully tuned, the model can produce unstable eigenvalues, leading to poor performance. Figure 8 illustrates this issue by showing the discrete-time eigenvalues of the standard and dissipative Koopman operators for the Van der Pol Oscillator System at a noise level of 0.0599, using a different weight initialization than in previous cases. For fair comparison, both standard and dissipative Koopman models are initialized with stable Koopman matrices. However, after 1,000 training epochs, some of the discrete-time eigenvalues in the Standard Koopman matrix move outside the unit circle, leading to a loss of stability while dissipative Koopman model never loses stability during training.
@@ V-B Learning Rate Scheduler
Figure 9b compares the performance of the model using constant learning rate (LR) scheduling and cyclic learning rate scheduling under different noise levels. The cyclic learning rate scheduling generally results in slightly lower mean errors and mean test errors compared to the constant learning rate. The time taken for training with cyclic learning rate scheduling is slightly higher due to the oscillation in learning rates.
@@ V-C Optimizer Switching
Figure 9c examines the effect of switching optimizers on the model’s performance. Different combinations of Adam and LBFGS optimizers were used, and their impact on mean errors, mean test errors, and training times was analyzed. Switching optimizers can lead to lower mean errors and test errors in some configurations, suggesting that optimizer switching can enhance training effectiveness. The combination of Adam and LBFGS every 500 epochs generally shows better performance with lower errors but increases the training time significantly. The mean error and test error for configurations where optimizers are switched show improved stability and reduced overfitting.
V-D Progressive roll-out length versus Constant roll-out length
Figure 9d compares the performance of the model using progressive roll-out length scheduling and constant roll-out length scheduling under different noise levels. The progressive roll-out scheduling shows lower mean errors and mean test errors compared to the constant roll-out length, indicating better learning and generalization. The time taken for training with progressive roll-out scheduling is higher due to the incremental increase in roll-out length.
@@ V-D Progressive roll-out length versus Constant roll-out length
Figure 9: Ablation study on the Van der Pol Oscillator system. (a) Effect of stabilized formulation of the Koopman operator. (b) Effect of learning rate scheduler. (c) Effect of optimizer switching. (d) Effect of progressive roll-out.
@@ VI Conclusions
In this work, we proposed a novel data-driven learning framework for the Koopman operator of nonlinear dynamical systems. Unlike existing frameworks that rely on ad-hoc observables or complex neural networks to approximate the Koopman invariant subspace, our approach utilizes observables chosen from known system dynamics with closure through the concept of Polyflow. This physics-informed basis significantly enhances the performance and robustness of the resulting models compared to ad-hoc choices. To ensure noise robustness and long-term stability of the lifted system, we designed a stable parameterization of the Koopman operator and employed a roll-out recurrent loss. Our stabilized parameterization guarantees a stable system, which can be equally accurate or, in some cases, even more accurate than the standard Koopman operator. Additionally, we addressed the curse of dimensionality in sampling from high-dimensional phase space by iteratively augmenting data from closed-loop trajectories. This approach not only improved performance in seen closed-loop scenarios but also enhanced the model’s ability to handle unseen control objectives effectively. We conducted an ablation study to demonstrate the effectiveness of several key components: switching to the LBFGS optimizer for fine-tuning, employing cyclic learning rate scheduling, utilizing progressive roll-out lengths for enhanced training outcomes, and incorporating a stabilized Koopman operator to ensure guaranteed system stability. Our numerical results on the Van der Pol Oscillator and the CartPole system demonstrated the superior performance of the proposed framework in prediction and model predictive control (MPC) of nonlinear systems, compared to other popular DMD-based methods. Overall, our framework offers a noise-robust approach for learning stable Koopman operators, providing a solid foundation for future research (e.g., Koopman for reinforcement learning [11]) and applications in nonlinear dynamical system control.